OpenStudy (anonymous):

Can you simplify 24x3/25y2*10y5/18x step by step for me.

OpenStudy (kc_kennylau):

\[\large\frac{24x^3}{25y^2}\cdot\frac{10y^5}{18x}\]Do you mean this? :)

OpenStudy (anonymous):

Yes :)

OpenStudy (anonymous):

Im studying for my math final... Im pretty nervous.

OpenStudy (kc_kennylau):